OpenStudy (anonymous):

Solve the quadratic equation. x^2 - 4x - 3 = 0

OpenStudy (anonymous):

(x-4)x=3 x^2=4x+3 (x-2)^2-7=0 x=2 x=2

OpenStudy (anonymous):

kumar that is wrong.

OpenStudy (anonymous):

How is that wrong

OpenStudy (anonymous):

since you cant factor that trinomial, youll have to plug it into the quadratic equation.

OpenStudy (anonymous):

I think I am right.

OpenStudy (anonymous):

the answer is 2\[\pm \sqrt{7}\]

OpenStudy (anonymous):

2 plus/minus root 7

OpenStudy (anonymous):

x=2+ sqroot 7

OpenStudy (anonymous):

which one should i pick??? (A) 2+- sq rt 7 (B) 4+-2 SQ RT 7 (c) 2+- sq rt 7 over 2 (D) 2+- SQ RT 28 OVER 2

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

kumar, anytime you take the square root of both sides of the equation the answer will be +/-.

OpenStudy (anonymous):

thax

OpenStudy (anonymous):

sq rt x-3 then outside the sq rt box is +6=2

OpenStudy (anonymous):

\[\sqrt{x-3} +6=2\]